# On a generalisation of Cameron’s base size conjecture

### v2

@@ 7. Preliminaries on exceptional groups
We will denote the number of elements of order $r$ in some group $H$ by $i_{r}(H)$. We will often denote the set of prime numbers at most $r$ by $\pi_{r}$ and the number of odd primes at most $r$ by $\pi^{\prime}_{r}$.
For the remainder of the paper we will write $\mathcal{M}_{G}$ for the set of core-free maximal subgroups of $G$ and $\mathcal{P}_{G}$ for the set of parabolic maximal subgroups of $G$, and finally, following the notation in [51], we will often write $x=s,u,\phi,\tau,\phi\tau$ to mean that $x$ is semisimple, unipotent, a field, graph, or graph-field automorphism respectively.
@@ 7.1. Subgroup structure
  • $\mathbf{G_{0}={}^{2}B_{2}(q):}$ In this case, Suzuki determines the maximal subgroups of $G$ up to conjugacy in [74] and Bray, Hold and Roney-Dougal record them in [12, Table 8.16].
  • $\mathbf{G_{0}={}^{2}G_{2}(q)^{\prime}\textbf{ or }{}^{3}D_{4}(q):}$ Here the subgroups in $\mathcal{M}_{G}$ up to conjugacy are determined by Kleidman in [48] and [50] respectively, and unless $q=3$, they are listed in [12, Tables 8.43, 8.51].
  • $\mathbf{G_{0}={}^{2}F_{4}(q)^{\prime}:}$ In the case $q\geqslant 8$ the maximal subgroups of $G$ are determined by Malle in [66] apart from the omission of three classes of maximal subgroups isomorphic to ${\rm PGL}_{2}(13)$ when $q=8$. This is straightened out in [33, Remark 4.11].
    If $q=2$, then the subgroups in $\mathcal{M}_{G}$ are determined up to conjugacy by Wilson in [76] with the exception of the omission of a unique class of maximal subgroups isomorphic to ${\rm U}_{3}(2).2$.
  • $\mathbf{G_{0}=G_{2}(q)^{\prime}:}$ Here the members of $\mathcal{M}_{G}$ are determined up to conjugacy in [32] for even $q$ and [48] for odd $q$. For $q=2^{f}$ and $f>1$ one can find them listed [12, Table 8.30] and for odd $q$ they are listed in [12, Tables 8.41, 8.42].
  • $\mathbf{G_{0}=F_{4}(q):}\textbf{ or }\mathbf{{E}_{6}^{\epsilon}(q):}$ The definitive source for those groups is [33], where Craven determines their maximal subgroups up to conjugacy. The members of $\mathcal{C}$ are listed in [33, Table 7, 8] and the members of $\mathcal{S}$ are listed in [33, Tables 1-3].
  • $\mathbf{G_{0}=E_{7}(q):}$ The members of $\mathcal{M}_{G}$ are discussed in [34] in this case. In particular, the members of $\mathcal{C}$ are listed in [34, Table 4.1] up to conjugacy, and as stated in [34, Theorem ], the collection $\mathcal{S}$ consists of the groups listed in [34, Table 1.1] up to conjugacy possibly together with some additional subgroups with socle ${\rm L}_{2}(q^{\prime})$ for $q\in\{7,8,9,13\}$.
  • $\mathbf{G_{0}=E_{8}(q):}$ The subgroups in $\mathcal{C}$ have been completely determined up to conjugacy. In particular, the parabolic and maximal rank subgroups of the form $N_{G}(\bar{H}_{\sigma})$ are listed in [56, Tables 5.1 and 5.2], the subfield subgroups and exotic local subgroups are listed in [57, Table 1], and finally for the non-maximal rank subgroups of the form $N_{G}(\bar{H}_{\sigma})$, one can find the possibilities for ${\rm soc}(\bar{H}_{\sigma})$ listed in [31, Table 3]. If $p>5$ then $\mathcal{C}$ also contains the Borovik subgroup.
    It is currently an open problem to determine the members of $\mathcal{S}$ even up to isomorphism. However, significant progress has been made towards a classification. The current state of the art regarding the groups in class $\mathcal{S}$ will not be relevant to us, but we refer the reader to [24, p.19] for more information.
@@ 7.2. Conjugacy classes
$\widehat{Q}(G,\tau)\leqslant\sum_{x\in\pi(|G|)}f(x,q),$
where
@@
When $H$ is non-parabolic, our standard approach will be to obtain sufficient bounds by inspecting the proofs of [25] and [19] when $H$. In certain rare cases, we will also need to improve existing bounds in the literature. For example, if $G_{0}=F_{4}(2)$ and $|H|>2^{22}$, then we will use the Character Table library in GAP to compute the precise value of ${\rm fpr}(x,G/H)$ for all $x\in G$ More details can be found in the proof of Proposition 8.4.
If $G$ has socle in $\mathcal{E}_{2}$, then obtaining fixed point ratio bounds that hold for all non-parabolic subgroups of $G$ is slightly more involved than in other cases, due to the larger number of possibilities for $H$, and so we carefully explain how to do this in Section 8. Surprisingly, if $G_{0}=E_{8}(q)$, then obtaining sufficient bounds is not as troublesome and can be done by careful inspection of [25]. These bounds may be of independent interest, so we record them in Tables 12, 13, 14, 15, and 16 for the reader’s convenience.
Our approach when $H$ is parabolic is rather different and we give detailed information on how we obtain bounds on ${\rm fpr}(x,G/H)$ in Section 9. If $G_{0}\in\mathcal{E}_{1}$, then the bounds in [51, Theorems 1–2] are always sufficient for unipotent and semisimple elements, and better bounds can be obtained by inspecting the proof of [51, 6.1] for field and graph-field automorphisms, if necessary. This is however not true if $G_{0}\in\mathcal{E}_{2}\cup\mathcal{E}_{3}$, where we will heavily use [51, 3.2] and [64] to derive better estimates for ${\rm fpr}(x,G/H)$ that are sensitive to ${\rm dim}\,x^{\bar{G}}$ for a given prime order element $x\in G$. For a unipotent element $x\in G$, it is always possible to compute precise values for ${\rm fpr}(x,G/H)$ using Lübeck’s values of the Foulkes functions for $G_{0}$ [64]. Now suppose that $x\in G$ is semisimple. If $G_{0}\in\mathcal{E}_{3}$, then an upper bound on the formula given by [51, 3.2] will be sufficient for our purposes but if $G_{0}\in\mathcal{E}_{2}$, then we will need to compute the precise value of ${\rm fpr}(x,G/H)$ to obtain our result. For field and graph-field automorphisms we will always appeal to a statement extracted from the proof of [51, 6.1] (see Proposition 9.4). Finally, for graph automorphisms when $G_{0}=E_{6}^{\epsilon}(q)$, we derive appropriate bounds in Proposition 9.5 when $q$ is odd using the approach described in  [51, 6.4], which we record in Table 18, and for even $q$, we consult [25, Table 6] and for the reader’s convenience we record this information in Table 17. We give more details in Section 9.
@@ 8. Fixed point ratios for non-parabolic actions
for every $H\in\mathcal{M}_{G}\setminus\mathcal{P}_{G}$ and we record those values in Tables 12, 13, 14, 15, and 16. As mentioned above, $g(x,q)$ is readily obtained by inspecting the proofs of [25] in the majority of cases, and we carry out the necessary computations for the cases where more work is required.
@@
Case 1. $x$ is semisimple: Here, [51, Theorem 2] implies that ${\rm fpr}(x,G/H)\leqslant 2q^{-12}$, and we claim that ${\rm fpr}(x,G/H)\leqslant q^{-14}$ if ${\rm dim}\,x^{\bar{G}}\geqslant 66$. Let $L$ denote the simply connected group of type $E_{6}^{\epsilon}(q)$ and recall that $L=G_{0}.Z$, where $Z=Z(L)$. Also recall that $Z\cong C_{3}$ if $q\equiv 1\allowbreak\mkern 4.0mu({\operator@font mod}\,\,3)$ and $Z=1$ otherwise.
@@
$|C_{L}(x)|\leqslant(q^{2}-1)(q^{3}+1)(q^{2}-q+1)(q^{2}+q+1)$
when $\epsilon=+$ and ${\rm dim}\,x^{\bar{G}}\geqslant 66$. Since $|x^{L}|$ can split into at most $G_{0}$-classes, we have
$|x^{G}|$ $\geqslant\frac{|E_{6}(q)|}{3q^{3}(q^{3}+1)(q^{2}+q+1)(q^{2}-q+1)}$
$\geqslant\frac{1}{3}q^{39}(q^{2}-1)(q^{3}-1)(q^{5}-1)(q^{6}+1)(q^{8}-1)(q^{9}-1)$
$\geqslant\frac{q^{39}(q^{2}-1)(q^{3}-1)(q^{5}-1)(q^{6}+1)(q^{8}-1)(q^{9}-1)}{3(q^{2}-1)}$
@@
If $|H|\leqslant q^{48}$, then we deduce that ${\rm fpr}(x,G/H)<2q^{-15}<q^{-14}$, as required, so now suppose $|H|>q^{48}$. Here, [33] tells us that $H$ must be of type $F_{4}$, in which case we verify that ${\rm fpr}(x,G/H)<q^{-14}$ by inspecting the proof of [25, 4.18].
Case 2. $x$ is unipotent: Here, $x\in G_{0}$. If $p>2$, then we check that the relevant bounds hold by inspecting [25, 4.14–4.20], so let us assume that $p=2$. If ${\rm dim}\,x^{\bar{G}}<40$, then the claim follows directly by [51, Theorem 2], so we may now assume that ${\rm dim}\,x^{\bar{G}}\geqslant 40$. That is, $x$ is in the $G$-class $A_{1}^{3}$. If $|H|\leqslant q^{32}$, or $H$ is of type $E_{6}^{\epsilon}(q^{1/2})$, then the claim is verified in the proofs of [25, 4.14, 4.20] respectively, so we may assume that $H=N_{G}(\bar{M}_{\sigma})$ where $\bar{M}$ is a $\sigma$-stable subgroup of $\bar{G}$.
First suppose that $\bar{M}=A_{1}A_{5}$. By Bala-Carter, the $\bar{M}$-class of $x$ corresponds to a pair $(L,P_{L^{\prime}})$, where $L$ is a Levi subgroup of $\bar{M}$ and $P_{L^{\prime}}$ is a distinguished parabolic subgroup of $L^{\prime}$. If $L$ is also a Levi subgroup of $\bar{G}$, then the $\bar{G}$-class has the same label as the $\bar{M}$-class, whilst if $L=A_{1}^{4}$, then we deduce that $x$ lies in the $G$-class $A_{1}^{3}$ via [54, Table 17]. Now it is shown in [25, 4.15] that ${\rm fpr}(x,G/H)<4q^{-\delta(x)}$, where $\delta(x)={\rm dim}\,x^{\bar{G}}-{\rm dim}\,(x^{\bar{G}}\cap\bar{M})$, and so we compute that ${\rm fpr}(x,G/H)<4q^{-20}$.
Next consider the case $\bar{M}=D_{5}T_{1}$. Here [33, Tables 9 and 10] give $H_{0}=H\cap G_{0}={\rm P\Omega}_{10}^{\epsilon}(q)\times\frac{q-\epsilon}{e}$, where $e=(3,q-\epsilon)$. Therefore, [51, 1.3] implies that
@@
Now suppose that $\bar{M}=T_{2}D_{4}.S_{3}$. Here we deduce via [33, Tables 9 and 10] that $x^{G}\cap H\subseteq\widetilde{H}$, where $\widetilde{H}={\rm O}_{8}^{+}(q)$ or $^{3}D_{4}(q)$. Hence, we have
@@
Finally, if $\bar{M}=F_{4}$, then we deduce via [53, Table A] that there are no involutions $x\in H$ with ${\rm dim}\,x^{\bar{G}}\geqslant 40$.
Case 3. $x$ is a graph or field or graph-field automorphism: If $r$ is odd and $|H|\leqslant q^{32}$ we observe that ${\rm fpr}(x,G/H)\leqslant q^{32-78(1-1/r)}$, as
@@
In all other cases the bounds derived in the proofs of [25, 4.14–4.20] are again sufficient, so the proof is complete. ∎
@@
If $H$ is almost simple with socle ${\rm Fi}_{22}$, then the fusion of conjugacy classes of $H$ in $G$ is computable in GAP using the Character Table Library, so we can compute precise fixed point ratios for all $x\in G$, as noted in Section 2.2. The same is true if $G={}^{2}E_{6}(2).2$ and $H={\rm SO}_{7}(3)$. So, in view of [25, 4.13], we may assume from now on that $|H|\leqslant 2^{32}$ or $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$.
Case 1. $x$ semisimple: If ${\rm dim}\,x^{\bar{G}}$, then the claim follows directly by [51, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}\geqslant 42$. If $H=N_{\bar{G}}(\bar{M}_{\sigma})$ with $\bar{M}=F_{4}$, then we note that $G\leqslant G_{0}.2$, and the character tables of $G_{0}$ and $H_{0}=H\cap G_{0}$ are both available in GAP. Since $x\in G_{0}$, we can compute the fusion of semisimple $H_{0}$-classes in $G_{0}$ and verify that the bounds in Table 13 are satisfied for all semisimple elements.
Next, if $H=N_{\bar{G}}(\bar{M}_{\sigma})$ with $\bar{M}=D_{5}T_{1}$ then the bounds in the proof of [25, 4.16] suffice.
In all remaining cases, we can infer from [33, Tables 9 and 10] that $|x^{G}\cap H|\leqslant 2^{38}$ and we immediately get ${\rm fpr}(x,G/H)\leqslant 2^{38}/|x^{G}|$ when ${\rm dim}\,x^{\bar{G}}>42$, so finally let us assume that ${\rm dim}\,x^{\bar{G}}=42$, so $C_{\bar{G}}(x)^{0}=T_{1}A_{5}$.
First suppose that $|H|\leqslant 2^{32}$. We can infer from [25, Table 7] that $|x^{G}|>2^{41}$, so we have ${\rm fpr}(x,G/H)<2^{-9}$.
Now assume that $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$. If $\bar{M}=A_{1}A_{5}$, then ${\rm fpr}(x,G/H)<2^{-14}$ according to [25, 4.15].
Finally, if $\bar{M}=T_{2}D_{4}.S_{3}$, then inspection of [33, Tables 9 and 10] gives
@@
Case 2. $x$ is unipotent: If $x$ is in class $A_{1}$, then [51, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant 2^{-5}$, so next assume that $x$ is in class $A_{1}^{2}$ or $A_{1}^{3}$. If $|H|\leqslant 2^{32}$, then the claim is proved in [25, 4.14], so we may now assume that $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$.
If $\bar{M}=A_{1}A_{5}$, then the bound ${\rm fpr}(x,G/H)\leqslant 4\cdot 2^{-\delta(x)}$ is still valid, as highlighted in [25, 4.15], and we find that $\delta(x)=16$ if $x$ is in $A_{1}^{2}$ and $\delta(x)=20$ if $x$ is in $A_{1}^{3}$, so the claim holds.
@@
Case 3. x is an involutory graph automorphism: Here, the bounds in Table 13 can be verified by inspecting [25, 4.14–4.20], so the proof is complete.
@@
Appealing to [25, 4.6] we find that either $|H|\leqslant q^{46}$, or $H$ is of type $E_{7}(q^{1/2})$, or $H=N_{G}(\bar{M}_{\sigma})$, where $\bar{M}=T_{1}E_{6}.2,A_{1}D_{6},A_{7}.2$, or $A_{1}F_{4}$.
Case 1. $x$ is semisimple: We first consider the case $q\geqslant 3$, as $q=2$ needs special attention. Here, [51, Theorem 2] implies that ${\rm fpr}(x,G/H)<q^{-19}$, so the desired bound holds if ${\rm dim}\,x^{\bar{G}}\leqslant 70$. We may now thus assume that ${\rm dim}\,x^{\bar{G}}>70$. Recall that if $L$ denotes the simply connected group of type $E_{7}(q)$, then $L=G_{0}.Z$, where $Z=Z(L)$ and $Z\cong C_{2}$ if $q$ is odd and $Z=1$ if $q$ is even. By carefully inspecting [39], we find that
$C_{L}(x)\leqslant q^{21}(q+1)(q^{2}-1)(q^{3}+1)(q^{4}-1)(q^{5}+1)(q^{6}-1)(q^{7}+1)$

### v3

@@ 7. Exceptional group preliminaries
We will denote the number of elements of order $r$ in some group $H$ by $i_{r}(H)$. We will often denote the set of prime numbers at most $r$ by $\pi_{r}$ and the set of odd primes at most $r$ by $\pi^{\prime}_{r}$.
Finally, for the remainder of the paper we will write $\mathcal{M}_{G}$ for the set of core-free maximal subgroups of $G$ and $\mathcal{P}_{G}$ for the set of parabolic maximal subgroups of $G$.
@@ 7.1. Subgroup structure
  • $\mathbf{G_{0}={}^{2}B_{2}(q):}$ In this case, Suzuki determines the maximal subgroups of $G$ up to conjugacy in [77], and Bray, Hold and Roney-Dougal record them in [12, Table 8.16].
  • $\mathbf{G_{0}={}^{2}G_{2}(q)^{\prime}\textbf{ or }{}^{3}D_{4}(q):}$ Here, the subgroups in $\mathcal{M}_{G}$ up to conjugacy are determined by Kleidman in [49] and [51] respectively, and unless $q=3$, they are listed in [12, Tables 8.43, 8.51].
  • $\mathbf{G_{0}={}^{2}F_{4}(q)^{\prime}:}$ In the case $q\geqslant 8$ the maximal subgroups of $G$ are determined by Malle in [67], apart from the omission of three classes of maximal subgroups isomorphic to ${\rm PGL}_{2}(13)$ when $q=8$. This is straightened out in [34, Remark 4.11].
    If $q=2$, then the subgroups in $\mathcal{M}_{G}$ are determined up to conjugacy by Wilson in [79], with the exception of the omission of a unique class of maximal subgroups isomorphic to ${\rm U}_{3}(2).2$.
  • $\mathbf{G_{0}=G_{2}(q)^{\prime}:}$ Here, the members of $\mathcal{M}_{G}$ are determined up to conjugacy in [33] for even $q$ and in [49] for odd $q$. For $q=2^{f}$ and $f>1$ one can find them listed [12, Table 8.30] and for odd $q$ they are listed in [12, Tables 8.41, 8.42].
  • $\mathbf{G_{0}=F_{4}(q):}\textbf{ or }\mathbf{{E}_{6}^{\epsilon}(q):}$ The definitive source for those groups is [34], where Craven determines their maximal subgroups up to conjugacy. The members of $\mathcal{C}$ are listed in [34, Tables 7-10] and the members of $\mathcal{S}$ are listed in [34, Tables 1-3].
  • $\mathbf{G_{0}=E_{7}(q):}$ The members of $\mathcal{M}_{G}$ are discussed in [35] in this case. In particular, the members of $\mathcal{C}$ are listed in [35, Table 4.1] up to conjugacy, and as stated in [35, 1.1], the collection $\mathcal{S}$ consists of the groups listed in [35, Table 1.1] up to conjugacy possibly together with some additional subgroups with socle ${\rm L}_{2}(q^{\prime})$ for $q\in\{7,8,9,13\}$.
  • $\mathbf{G_{0}=E_{8}(q):}$ The subgroups in $\mathcal{C}$ have been completely determined up to conjugacy. In particular, the parabolic and maximal rank subgroups of the form $N_{G}(\bar{H}_{\sigma})$ are listed in [57, Tables 5.1 and 5.2], the subfield subgroups and exotic local subgroups are listed in [58, Table 1], and finally for the non-maximal rank subgroups of the form $N_{G}(\bar{H}_{\sigma})$, one can find the possibilities for ${\rm soc}(\bar{H}_{\sigma})$ listed in [32, Table 3]. If $p>5$ then $\mathcal{C}$ also contains the Borovik subgroup.
    It is currently an open problem to determine the members of $\mathcal{S}$ even up to isomorphism. However, significant progress has been made towards a classification. The current state of the art regarding the groups in class $\mathcal{S}$ will not be relevant to us, but we refer the reader to [25, p.19] for more information.
@@ 7.2. Conjugacy classes
$\widehat{Q}(G,\tau)\leqslant\sum_{x\in\pi}f(x,q),$
where $\pi$ denotes the set of prime order elements in $G$ and
@@
When $H$ is non-parabolic, our standard approach will be to obtain sufficient bounds by inspecting the proofs of [26] and [19]. In certain rare cases, we will also need to improve existing bounds in the literature. For example, if $G_{0}=F_{4}(2)$ and $|H|>2^{22}$, then we will use the Character Table library in GAP to compute the precise value of ${\rm fpr}(x,G/H)$ for all $x\in G$ More details can be found in the proof of Proposition 8.4.
If $G$ has socle in $\mathcal{E}_{2}$, then obtaining fixed point ratio bounds that hold for all non-parabolic subgroups of $G$ is slightly more involved than in other cases, due to the larger number of possibilities for $H$. We carefully explain how to do this in Section 8. These bounds may be of independent interest, so we record them in Tables 13, 14, 15, 16, and 17 for the reader’s convenience. Surprisingly, if $G_{0}=E_{8}(q)$, then obtaining sufficient bounds is not as troublesome and can be done by careful inspection of [26].
Our approach when $H$ is parabolic is rather different and we give detailed information on how we obtain bounds on ${\rm fpr}(x,G/H)$ in Section 9. If $G_{0}\in\mathcal{E}_{1}$, then the bounds in [52, Theorems 1–2] are always sufficient for unipotent and semisimple elements, and better bounds can be obtained by inspecting the proof of [52, 6.1] for field and graph-field automorphisms, if necessary. This is however not true if $G_{0}\in\mathcal{E}_{2}\cup\mathcal{E}_{3}$, where we will heavily use [52, 3.2] and [65] to derive better estimates for ${\rm fpr}(x,G/H)$ that are sensitive to ${\rm dim}\,x^{\bar{G}}$ for a given prime order element $x\in G$. For a unipotent element $x\in G$, it is always possible to compute precise values for ${\rm fpr}(x,G/H)$ using Lübeck’s values of the Foulkes functions for $G_{0}$ [65].
Now suppose that $x\in G$ is semisimple. If $G_{0}\in\mathcal{E}_{3}$, then an upper bound on the formula given by [52, 3.2] will be sufficient for our purposes but if $G_{0}\in\mathcal{E}_{2}$, then we will need to compute the precise value of ${\rm fpr}(x,G/H)$ to obtain our result. For field and graph-field automorphisms we will always appeal to a statement extracted from the proof of [52, 6.1] (see Proposition 9.4). Finally, for graph automorphisms when $G_{0}=E_{6}^{\epsilon}(q)$, we derive appropriate bounds in Proposition 9.5 when $q$ is odd using the approach described in  [52, 6.4], which we record in Table 20, and for even $q$, we consult [26, Table 7] and for the reader’s convenience we record this information in Table 19. We give more details in Section 9.
@@ 8. Fixed point ratios for non-parabolic actions
for every $H\in\mathcal{M}_{G}\setminus\mathcal{P}_{G}$ and we record those values in Tables 13, 14, 15, 16, and 17. As mentioned above, $g(x,q)$ is readily obtained by inspecting the proofs of [26] in the majority of cases, and we carry out the necessary computations here for the cases where more work is required.
@@
Case 1. $x$ is semisimple: Here, [52, Theorem 2] implies that ${\rm fpr}(x,G/H)\leqslant 2q^{-12}$, and we claim that ${\rm fpr}(x,G/H)\leqslant q^{-14}$ if ${\rm dim}\,x^{\bar{G}}\geqslant 66$. Let $L$ denote the simply connected group of type $E_{6}^{\epsilon}(q)$ and recall that $L=G_{0}.Z$, where $Z=Z(L)$. Also recall that $Z\cong C_{3}$ if $q\equiv 1\allowbreak\mkern 4.0mu({\operator@font mod}\,\,3)$ and $Z=1$ otherwise.
@@
$|C_{L}(x)|\leqslant q^{3}(q^{2}-1)(q^{3}+1)(q^{2}-q+1)(q^{2}+q+1)$
when $\epsilon=+$ and ${\rm dim}\,x^{\bar{G}}\geqslant 66$. Since $|x^{L}|$ can split into at most 33 $G_{0}$-classes, we have
$|x^{G}|$ $\geqslant\frac{|E_{6}(q)|}{3q^{3}(q^{2}-1)(q^{3}+1)(q^{2}+q+1)(q^{2}-q+1)}$
$\geqslant\frac{1}{3}q^{33}(q^{2}-1)(q^{3}-1)(q^{5}-1)(q^{6}+1)(q^{8}-1)(q^{9}-1)$
$\geqslant\frac{q^{39}(q^{2}-1)(q^{3}-1)(q^{5}-1)(q^{8}-1)(q^{9}-1)}{3(q^{2}-1)}$
@@
If $|H|\leqslant q^{48}$, then we deduce that ${\rm fpr}(x,G/H)<2q^{-15}<q^{-14}$, as required, so now suppose $|H|>q^{48}$. Here, [34] tells us that $H$ must be of type $F_{4}$, in which case we verify that ${\rm fpr}(x,G/H)<q^{-14}$ by inspecting the proof of [26, 4.19].
Case 2. $x$ is unipotent: Here, $x\in G_{0}$. If $p>2$, then we check that the relevant bounds hold by inspecting [26, 4.15–4.21], so let us assume that $p=2$. If ${\rm dim}\,x^{\bar{G}}<40$, then the claim follows directly by [52, Theorem 2], so we may now assume that ${\rm dim}\,x^{\bar{G}}\geqslant 40$. That is, $x$ is in the $G$-class $A_{1}^{3}$. If $|H|\leqslant q^{32}$, or $H$ is of type $E_{6}^{\epsilon}(q^{1/2})$, then the claim is verified in the proofs of [26, 4.15, 4.21] respectively, so we may assume that $H=N_{G}(\bar{M}_{\sigma})$ where $\bar{M}$ is a $\sigma$-stable subgroup of $\bar{G}$.
First suppose that $\bar{M}=A_{1}A_{5}$. By Bala-Carter, the $\bar{M}$-class of $x$ corresponds to a pair $(L,P_{L^{\prime}})$, where $L$ is a Levi subgroup of $\bar{M}$ and $P_{L^{\prime}}$ is a distinguished parabolic subgroup of $L^{\prime}$. If $L$ is also a Levi subgroup of $\bar{G}$, then the $\bar{G}$-class has the same label as the $\bar{M}$-class, whilst if $L=A_{1}^{4}$, then we deduce that $x$ lies in the $G$-class $A_{1}^{3}$ via [55, Table 17]. Now it is shown in [26, 4.16] that ${\rm fpr}(x,G/H)<4q^{-\delta(x)}$, where $\delta(x)={\rm dim}\,x^{\bar{G}}-{\rm dim}\,(x^{\bar{G}}\cap\bar{M})$, and so we compute that ${\rm fpr}(x,G/H)<4q^{-20}$.
Next consider the case $\bar{M}=D_{5}T_{1}$. Here [34, Tables 9 and 10] give $H_{0}=H\cap G_{0}={\rm P\Omega}_{10}^{\epsilon}(q)\times\frac{q-\epsilon}{e}$, where $e=(3,q-\epsilon)$. Therefore, [52, 1.3] implies that
@@
Now suppose that $\bar{M}=T_{2}D_{4}.S_{3}$. Here we deduce via [34, Tables 9 and 10] that $x^{G}\cap H\subseteq\widetilde{H}$, where $\widetilde{H}={\rm O}_{8}^{+}(q)$ or $^{3}D_{4}(q)$. Hence, we have
@@
Finally, if $\bar{M}=F_{4}$, then we deduce via [54, Table A] that there are no involutions $x\in H$ with ${\rm dim}\,x^{\bar{G}}\geqslant 40$.
Case 3. $x$ is a graph or field or graph-field automorphism: If $r$ is odd and $|H|\leqslant q^{32}$ we observe that ${\rm fpr}(x,G/H)\leqslant 6q^{32-78(1-1/r)}$, as
@@
In all other cases the bounds derived in the proofs of [26, 4.15–4.21] are again sufficient, so the proof is complete. ∎
@@
If $H$ is almost simple with socle ${\rm Fi}_{22}$, then the fusion of conjugacy classes of $H$ in $G$ is computable in GAP using the Character Table Library, so we can compute precise fixed point ratios for all $x\in G$, as noted in Section 2.2. The same is true if $G={}^{2}E_{6}(2).2$ and $H={\rm SO}_{7}(3)$. So, in view of [26, 4.14], we may assume from now on that $|H|\leqslant 2^{32}$ or $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$.
Case 1. $x$ semisimple: If ${\rm dim}\,x^{\bar{G}}<42$, then the claim follows directly by [52, Theorem 2], so we may assume that ${\rm dim}\,x^{\bar{G}}\geqslant 42$. If $H=N_{\bar{G}}(\bar{M}_{\sigma})$ with $\bar{M}=F_{4}$, then we note that $G\leqslant G_{0}.2$, and the character tables of $G_{0}$ and $H_{0}=H\cap G_{0}$ are both available in GAP. Since $x\in G_{0}$, we can compute the fusion of semisimple $H_{0}$-classes in $G_{0}$ and verify that the bounds in Table 14 are satisfied for all semisimple elements.
Next, if $H=N_{\bar{G}}(\bar{M}_{\sigma})$ with $\bar{M}=D_{5}T_{1}$ then the bounds in the proof of [26, 4.17] suffice.
In all remaining cases, we can infer from [34, Tables 9 and 10] that $|x^{G}\cap H|\leqslant 2^{38}$ and we immediately get ${\rm fpr}(x,G/H)\leqslant 2^{38}/|x^{G}|$ when ${\rm dim}\,x^{\bar{G}}>42$, so finally let us assume that ${\rm dim}\,x^{\bar{G}}=42$, so $C_{\bar{G}}(x)^{0}=T_{1}A_{5}$.
First suppose that $|H|\leqslant 2^{32}$. We can infer from [26, Table 9] that $|x^{G}|>2^{41}$, so we have ${\rm fpr}(x,G/H)<2^{-9}$.
Now assume that $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$. If $\bar{M}=A_{1}A_{5}$, then ${\rm fpr}(x,G/H)<2^{-14}$ as shown in [26, 4.16].
Finally, if $\bar{M}=T_{2}D_{4}.S_{3}$, then inspection of [34, Tables 9, 10] gives
@@
Case 2. $x$ is unipotent: If $x$ is in class $A_{1}$, then [52, Theorem 2] gives ${\rm fpr}(x,G/H)\leqslant 2^{-5}$, so next assume that $x$ is in class $A_{1}^{2}$ or $A_{1}^{3}$. If $|H|\leqslant 2^{32}$, then the claim is proved in [26, 4.15], so we may now assume that $H=N_{\bar{G}}(\bar{M}_{\sigma})$ where $\bar{M}$ is some $\sigma$-stable subgroup of $\bar{G}$.
If $\bar{M}=A_{1}A_{5}$, then the bound ${\rm fpr}(x,G/H)\leqslant 4\cdot 2^{-\delta(x)}$ is still valid, as highlighted in [26, 4.16], and we find that $\delta(x)=16$ if $x$ is in $A_{1}^{2}$ and $\delta(x)=20$ if $x$ is in $A_{1}^{3}$, so the claim holds.
@@
Case 3. x is an involutory graph automorphism: Here, the bounds in Table 14 can be verified by inspecting [26, 4.15–4.21], so the proof is complete.
@@
Appealing to [26, 4.7] we find that either $|H|\leqslant q^{46}$, or $H$ is of type $E_{7}(q^{1/2})$, or $H=N_{G}(\bar{M}_{\sigma})$, where $\bar{M}=T_{1}E_{6}.2,A_{1}D_{6},A_{7}.2$, or $A_{1}F_{4}$.
Case 1. $x$ is semisimple: We first consider the case $q\geqslant 3$, as $q=2$ needs special attention. Here, [52, Theorem 2] implies that ${\rm fpr}(x,G/H)<q^{-19}$, so the desired bound holds if ${\rm dim}\,x^{\bar{G}}\leqslant 70$. We may now thus assume that ${\rm dim}\,x^{\bar{G}}>70$. Recall that if $L$ denotes the simply connected group of type $E_{7}(q)$, then $L=Z.G_{0}$, where $Z=Z(L)$ and $Z\cong C_{2}$ if $q$ is odd and $Z=1$ if $q$ is even. By carefully inspecting [40], we find that
$C_{L}(x)\leqslant|{\rm SU}_{7}(q)|(q+1)=q^{21}(q+1)(q^{2}-1)(q^{3}+1)(q^{4}-1)(q^{5}+1)(q^{6}-1)(q^{7}+1)$